A solid cylinder of mass 4kg and radius 40cm rotates about its axis with a frequency of 1/pi what is its rotational kinetic energy?










1
Expert's answer
2022-05-15T17:02:22-0400

"I=\frac{mr^2}2,"

"\omega=2\pi\nu=2,"

"E=\frac{I\omega^2}2=0.64~J."